1. Consider an array of n distinct values in which the first n − 1 values are sorted, and the last element is not. (It could be smaller than the first element, larger than element n − 1, or anywhere in between.) Give the worst-case number of comparisons that Insertion Sort will perform in this scenario. You can give your answer in terms of big-Theta if you wish to ignore low-order terms and constants.
Answer #1

The subarrays from [1,i] for 2\le i\le n-1 are already sorted as per the hypothesis. So the I-th element must be compared with all previous (i-1) elements. This itself gives rise to

\sum_{i=2}^{n-1}(i-1)=\frac{(n-2)(n-1)}{2} which is Θ(n2)

The last insertion may take 1 comparison or n-1 comparisons depending on how large the last element is

In any case, the number of comparisons is Θ(n2)